OpenStudy (anonymous):

A researcher studying the nutritional value of a new candy places a 6.40-gram sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increase is 2.32 °C. If the heat capacity of the calorimeter is 33.10 kJ·K–1, how many nutritional Calories are there per gram of the candy?

OpenStudy (aaronq):

look at the units of the heat capacity, their not specifying mass. so the typical "\(q=m*C_p\Delta T\)" formula is reduced to "\(q=C_p*\Delta T\)". So find q (heat energy), convert it to the specified units (calories) and divide it by the mass of the candy to get "calories per gram".

OpenStudy (anonymous):

Thanks, that worked! I just had to plug in the correct conversion factor. 2.32˚C(33.10kj•K-1)=76.79 76.79/4.184=18.35 18.35/6.4=2.87Cal/g